What is the pH of a solution containing 5.6 mg / L of KOH, assuming it is completely dissociated?
1
Expert's answer
2020-12-18T08:38:26-0500

concentration of KOH = 5.6mg/L = 5.6mg/dm³

5.6mg if KOH = 5.6×10-³/56 mol = 1 × 10-⁴ moles


Therefore,the concentration of KOH = 1×10-⁴moles/dm³ = 10-4M


Now as pH is molar concentration of hydrogen ions of a specie, so we need to know molar concentration of hydroxide ion


"KOH \to K^+ + OH^-"


Therefore the concentration of the hydroxide ions is 10-⁴M.


Using the formula pOH = -log[OH-] = -log(10-4) = 4


but pH + pOH = 14

pH = 14 - 4 = 10


Therefore the pH of the potassium hydroxyl is 10